OpenStudy (anonymous):

A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits. I'm not exactly sure what to do. Help Please?.

OpenStudy (paki):

0.01 is the accepted tolerance for the drill bit... @Princess_Mia

OpenStudy (amistre64):

do we agree that |.01| = .01 and |-.01| = .01?

OpenStudy (amistre64):

if so then solve for both cases: f(x) = k and f(x) = -k

OpenStudy (anonymous):

what does f and k stand for? @amistre64

OpenStudy (paki):

@Princess_Mia please participate in the question.....

OpenStudy (amistre64):

they are generalties. given that you have posted; |f(x)| = k

OpenStudy (anonymous):

ok @paki

OpenStudy (amistre64):

f(x) is read as: a function of x, and not f, times x

OpenStudy (anonymous):

ohh ok

OpenStudy (amistre64):

lost connections, sites too broken to use for me

OpenStudy (paki):

for solving the equation we will need to split the main equation for two... |x − 0.525| = 0.01.......... |x − 0.525| = +0.01 and |x − 0.525| = -0.01 + and - are for that small tolerance... rite...